OpenStudy (blackstreet23):

Express the speed limit of 65 miles/hour in terms m/s

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 65{\rm miles}{~/~}1{\rm hour} }\)

OpenStudy (solomonzelman):

convert hours to seconds. 1 hour = 60 minutes 1 minute = 60 seconds Therefore: 1 hour = 360 seconds

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 65{\rm miles}{~/~}360{\rm seconds} }\)

OpenStudy (solomonzelman):

then convert miles to meters.

OpenStudy (blackstreet23):

wouldnt it be 3600?

OpenStudy (solomonzelman):

1 mile = 1609.34 meters

OpenStudy (blackstreet23):

m/s means meters per second right?

OpenStudy (solomonzelman):

1609.34 × 65 = 104609.1 \(\large\color{black}{ \displaystyle 104609.1~{\rm miles}{~/~}360{\rm seconds} }\)

OpenStudy (solomonzelman):

divide top and bottom by 360, and you get?

OpenStudy (blackstreet23):

no i mean. 1 hour = 3600 seconds right? 1*60*60

OpenStudy (solomonzelman):

oh, true, my bad.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 104607.1~{\rm miles}{~/~}3600{\rm seconds} }\) then divide by 3600 on top and bottom

OpenStudy (solomonzelman):

http://www.wolframalpha.com/input/?i=%281609.34%C3%9765%29%2F3600 \(\large\color{black}{ \displaystyle 29.057527777777...~{\rm miles}{~/~}{\rm seconds} }\)

OpenStudy (blackstreet23):

I am not understanding

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 29.1{\rm m}{/}{\rm s} }\)

OpenStudy (solomonzelman):

I have this: \(\large\color{black}{ \displaystyle 104607.1~{\rm miles}{~/~}3600{\rm seconds} }\) then I am dividing by 3600 on top and bottom so that I get meters per seconds.

OpenStudy (solomonzelman):

and this is what I got: \(\large\color{black}{ \displaystyle 29.1{\rm m}{/}{\rm s} }\)

OpenStudy (blackstreet23):

could you show me the whole process that you are doing, please ?

OpenStudy (blackstreet23):

\[65 *\frac{ miles }{ hour } = 65 * \frac{ 1609.34 meters }{ 3600 seconds }\]

OpenStudy (blackstreet23):

is it?

OpenStudy (blackstreet23):

@freckles @SolomonZelman @amistre64

OpenStudy (blackstreet23):

29.0575277778

OpenStudy (blackstreet23):

ohh i see now

OpenStudy (blackstreet23):

Thanks a lot !